An organ pipe sounds two successive tones at 228.6 Hz and 274.3 Hz. Determine whether the pipe is open at both ends or open at one end and closed at the other.

a)The pipe is open at both ends.

b)The pipe is open at one end and closed at the other.

* Please explain clearly how you calculated your answer
Answer #1

(0 closeal or gon pipa Pipa raan of vibratuns is seeeni over tome hne Frequency n5= 51 -囪 - ni 2-е equencius f Harmmica areif 2p+-) x4S-7 = 228-6-0 [By@า ( 2 fptr) +) x 4 s . 7ニ274.3 → @p+3) x, s.)- 2)1.3 -e uaftm weof P Hence, o scillahn ame nut